OpenStudy (anonymous):

Alex is saving money because she would like to purchase the new video gaming system that everyone has been talking about. After two weeks of saving, she had $47. After 24 weeks of saving, she had $430. Approximately how many dollars per week did Alex save on average between weeks 2 and 24? $15.96 per week $17.41 per week $17.91 per week $19.88 per week

OpenStudy (matimaticas):

B

OpenStudy (anonymous):

24-2=22 weeks of saving 430-47=383 383/22=17.41

OpenStudy (anonymous):

That's what I got, Thank You @Venny

OpenStudy (anonymous):

sorrey... 430-47=383 24-2=22 divide 383 by 22...

OpenStudy (anonymous):

You need to subtract 47 first